OpenStudy (luigi0210):

Integrate:

OpenStudy (luigi0210):

\[\LARGE \int \sqrt{5+4x-x^2}~dx\]

OpenStudy (luigi0210):

@dan815

OpenStudy (dan815):

so complete square first of all

OpenStudy (dan815):

integral sqrt(9-(x-2)^2) u=(/x-2) du=dx

OpenStudy (dan815):

then you can do trig sub

OpenStudy (dan815):

integral sqrt (9-u^2) integral sqrt(9-9 sin^2 x) u=3sinx du= 3 cos dx sqrt(9cos^2) 3cos * 3*cos dx

OpenStudy (dan815):

so we end up with integral 9cos^2 dx

OpenStudy (dan815):

should be pretty straight forward from here. apply one of those hand angle formulas or something and solve it